OpenStudy (anonymous):

Simplify completely quantity 12 x plus 36 over quantity x squared minus 4 x minus 21 and find the restrictions on the variable.

OpenStudy (anonymous):

@amistre64 @Hero @Mandre

OpenStudy (phi):

can you factor the top and bottom ?

OpenStudy (anonymous):

I tried... I think it's wrong

OpenStudy (phi):

can you factor (12x+36) ?

OpenStudy (anonymous):

6(2x+6)

OpenStudy (phi):

yes, but notice you can go even further

OpenStudy (anonymous):

or 12 maybe

OpenStudy (anonymous):

12(x+3)

OpenStudy (phi):

yes, now the bottom x^2 -4x -21 first list all pairs of number that multiply to give you 21 1,21 3,7 are all of them the - on the -21 means they will have opposite signs the -4x means they add up to -4 which pair should you use ?

OpenStudy (anonymous):

3 and -7?

OpenStudy (phi):

so the factors are (x+3)(x-7) you have \[ \frac{12(x+3)}{(x-7)(x+3)} \] do you see what you can "cancel"?

OpenStudy (anonymous):

x+3

OpenStudy (anonymous):

so we get 12/x-7?

OpenStudy (anonymous):

how do we find the restrictions?

OpenStudy (phi):

yes , but use parens: 12/(x-7) now find the restrictions on the variable. you do not want to divide by zero. in the original you were dividing by (x+3)(x-7) so you should not let x be certain values or you will get a zero what should x NOT be ?

OpenStudy (anonymous):

7

OpenStudy (phi):

yes. because (x-7) with x= 7 gives (7-7) = zero and (x+3)(x-7) becomes (7+3)*0 = 0 and we do not want to divide by 0 but there is another x that makes the bottom zero.

OpenStudy (anonymous):

I don't know

OpenStudy (phi):

if x+3 is 0, the bottom becomes zero

OpenStudy (anonymous):

-3?

OpenStudy (phi):

yes. you could "solve" this by saying either x-7= 0 add 7 to both sides x-7 +7 = 0 +7 x +0 = 7 x=7 or x+3= 0 add -3 to both sides x+3-3= 0-3 x= -3 so exclude x=-3 and x=7, because they would cause divide by 0

OpenStudy (anonymous):

Okay, I think I get it now :D do you think you think you could help me with the trinomial factoring again, though? I'm still a little sketchy on that. Let's say I have x^2-15x+36, how would I factor that? I can't think of any numbers that multiply to 36 and add to -15...

OpenStudy (phi):

look at the +36 list all the pairs that multiply to give 36 it helps if you (1) know the multiplication table (2) know some "divisibility rules" see https://en.wikipedia.org/wiki/Divisibility_rule#Divisibility_rules_for_numbers_1.E2.80.9320 I learned the rules for 2,3,4,5,6 can you list all pairs for 36 ?

OpenStudy (anonymous):

1 36, 2 18, 3 12, 4 9, 6 6

OpenStudy (phi):

yes, now the + sign on the 36 means the factors have the same sign (- means the factors would have different signs) the -15x means the larger factor is minus (so that means both factors are minus) they add to -15 which pair, when both numbers are minus, add to -15 ?

OpenStudy (anonymous):

ohhhh okay, so it would be -12 and -3?

OpenStudy (phi):

so you have (x-12)(x-3) as the factors multiply out to check: x^2 -12x -3x +36 = x^2 -15x +36

OpenStudy (phi):

if you make notes on how to do it, do a bunch of examples, you will learn how to do it.

OpenStudy (anonymous):

okay, thanks!